### v2

# Static for Dynamic: Towards a Deeper Understanding of Dynamic Facial Expressions Using Static Expression Data

Abstract
Dynamic facial expression recognition (DFER) infers emotions from the temporal evolution of expressions, unlike static facial expression recognition (SFER), which relies solely on a single snapshot. This temporal analysis provides richer information and promises greater recognition capability. However, current DFER methods often exhibit unsatisfied performance largely due to fewer training samples compared to SFER. Given the inherent correlation between static and dynamic expressions, we hypothesize that leveraging the abundant SFER data can enhance DFER. To this end, we propose Static-for-Dynamic (S4D), a unified dual-modal learning framework that integrates SFER data as a complementary resource for DFER. Specifically, S4D employs dual-modal self-supervised pre-training on facial images and videos using a shared Vision Transformer (ViT) encoder-decoder architecture, yielding improved spatiotemporal representations. The pre-trained encoder is then fine-tuned on static and dynamic expression datasets in a multi-task learning setup to facilitate emotional information interaction. Unfortunately, vanilla multi-task learning in our study results in negative transfer. To address this, we propose an innovative Mixture of Adapter Experts (MoAE) module that facilitates task-specific knowledge acquisition while effectively extracting shared knowledge from both static and dynamic expression data. Extensive experiments demonstrate that S4D achieves a deeper understanding of DFER, setting new state-of-the-art performance on FERV39K, MAFW, and DFEW benchmarks, with weighted average recall (WAR) of 53.65%, 58.44%, and 76.68%, respectively. Additionally, a systematic correlation analysis between SFER and DFER tasks is presented, which further elucidates the potential benefits of leveraging SFER.
Index Terms:
Dynamic facial expression recognition, mixture of experts, self-supervised learning, vision transformer.
I Introduction
Facial expression recognition (FER) is essential in fields such as human-computer interaction [1], mental health diagnosis [2], and driving safety [3]. Traditional FER methods focus on static facial expression recognition (SFER), which captures single moments of expressions from images. However, static images cannot fully reflect the dynamic changes of emotions over time. In contrast, video-based data provides richer temporal information and offers a more comprehensive view of emotions, prompting a shift towards dynamic facial expression recognition (DFER) [4].
Refer to captionMAFWFERV39KDFEW
Figure 1: Performance comparison between previous SOTA methods [5, 6, 7, 8] and our proposed S4D on FERV39K [9], MAFW [10], and DFEW [11] datasets. Unweighted average recall (UAR, %) and weighted average recall (WAR, %) are reported. S4D, which incorporates static expression knowledge through a unified dual-modal learning framework, consistently outperforms the baseline method, VideoMAE [8], previously pre-trained on VoxCeleb2 [12], across all these real-world DFER datasets.
To support the development and evaluation of DFER algorithms, researchers have constructed various datasets, including lab-controlled datasets and in-the-wild datasets. Lab-controlled datasets, such as CK+ [13], MMI [14], and Oulu-CASIA [15], are collected under controlled laboratory environments and contain exaggerated expressions performed by participants under specific instructions. However, the applicability of these datasets in real-world scenarios is limited. To overcome this limitation, researchers have started building large-scale in-the-wild datasets, such as DFEW [11], FERV39K [9], and MAFW [10]. These datasets are collected from real-world scenarios, including movies, TV shows, and online videos, and encompass a wide range of head movements, illumination variations, and spontaneous expressions. The emergence of these in-the-wild datasets has provided valuable data supporting the development of robust and practical DFER algorithms.
Despite the richer temporal cues inherent in video clips, which theoretically enable superior recognition, in-the-wild DFER lags behind its SFER counterpart. This performance disparity stems primarily from the limited scale and diversity of DFER datasets. While SFER benefits from large, diverse datasets containing millions of images and labels, DFER datasets are relatively smaller and less varied [7]. However, the intrinsic link between these tasks presents a valuable opportunity to improve DFER performance by leveraging SFER data. Specifically, SFER images often capture moments of pronounced emotional intensities and inherently contain highly discriminative features crucial for understanding dynamic expressions. From a temporal perspective, these static images can be viewed as critical snapshots that capture essential moments within dynamic expression sequences. Furthermore, the shared categorical labels and common semantic space between SFER and DFER, as evidenced by the semantic similarity of their class representations (see Fig. 4), suggest that leveraging the extensive SFER data can significantly enhance DFER model training while reducing reliance on scarce, temporally annotated video data.
Motivated by the inherent connection between SFER and DFER, recent research has begun exploring various approaches to transfer static knowledge to the dynamic domain. Representative works include AEN [16] and S2D [7]. AEN [16] pioneered this direction by integrating multi-level semantic features extracted from SFER models and employing emotion-guided loss functions to enhance recognition accuracy. Subsequently, S2D [7] further advances the field by incorporating Temporal Modeling Adapters (TMAs) to extend pre-trained SFER models for DFER tasks. Although these approaches have shown promising results, they primarily focus on feature-level knowledge transfer using pre-trained SFER models, potentially limiting the full exploitation of static expression information. We argue that a more comprehensive approach, utilizing both data (i.e., images) and corresponding labels from static datasets, could better leverage SFER knowledge to enhance DFER performance.
Based on these observations, we propose a novel framework, Static-for-Dynamic (S4D), which enhances the understanding of dynamic facial expressions by fully leveraging static expression data at both the input- and label- levels. To the best of our knowledge, we represent the first systematic exploration and analysis of the inherent correlations between SFER and DFER tasks, integrating both modalities in terms of input data (i.e., facial images and videos) and emotion labels. The S4D framework employs a unified dual-modal learning approach to seamlessly merge SFER and DFER tasks, thereby advancing DFER performance. Specifically, S4D comprises two key stages: Dual-Modal Pre-Training and Joint Fine-Tuning. During the pre-training stage, we utilize a shared Vision Transformer (ViT) [17] encoder-decoder architecture combined with Masked Autoencoders (MAE) [18] to perform masked modeling on both image and video modalities. This approach allows the model to efficiently learn generalizable spatiotemporal representations from both static and dynamic facial data. In the fine-tuning stage, the pre-trained encoder is jointly optimized using both FER image and video datasets with a multi-task learning setup, facilitating cross-modal interaction of emotional information. However, as highlighted in previous works [19, 20], directly employing multi-task learning in related FER tasks may result in a negative transfer. We also observed similar performance degradation when applying straightforward multi-task learning to SFER and DFER tasks. To address this issue, we introduce a novel Mixture of Adapter Experts (MoAE) module, which operates in parallel with the original Feed-Forward Network (FFN) in the ViT layers. This innovative design enables the model to simultaneously capture task-agnostic knowledge through the FFN and task-specific features via the MoAE, thereby preventing negative transfer and enhancing emotional information interaction. By integrating the MoAE module within the unified dual-modal learning framework, the S4D framework effectively integrates complementary information from static facial expression data, leading to a deeper understanding of dynamic facial expressions and significant improvements in DFER performance. The code and model are publicly available here [^1] https://github.com/MSA-LMC/S4D.
We summarize our main contributions as follows:
  • $\bullet$
    S4D, a Unified Dual-Modal Learning Framework. This framework seamlessly integrates dual-modal pre-training and joint fine-tuning on FER images and videos. Such a comprehensive approach yields rich and powerful spatiotemporal representations during pre-training and significantly improves expression semantic understanding through joint fine-tuning, ultimately achieving superior DFER performance. Incidentally, our final single model can perform both SFER and DFER tasks simultaneously.
  • $\bullet$
    MoAE, a Mixture of Adapter Experts Module. We incorporate the MoAE module into the ViT layers of the S4D encoder, operating in parallel with the FFN. This design allows FFN to focus on task-agnostic knowledge, while MoAE captures task-specific knowledge, effectively alleviating the negative transfer between SFER and DFER tasks and enabling more discriminative feature learning.
  • $\bullet$
    Analysis of Task Correlation and SOTA Performance. We present a systematic correlation analysis between SFER and DFER tasks from the perspectives of semantic and expert pathways, providing valuable insights into their inherent characteristics. Additionally, as shown in Fig. 1, our approach achieves substantial improvements over existing state-of-the-art methods in unweighted average recall (UAR) and weighted average recall (WAR), particularly surpassing the VideoMAE [8] baseline.
II Related Work
Figure 2: Overview of our proposed S4D framework. We utilize Vision Transformer (ViT) [17] as the backbone and pre-train it on facial image and video datasets using Masked Autoencoders [21]. The pre-trained ViT encoder is then used to initialize the S4D encoder, which is further fine-tuned on static and dynamic FER datasets. The proposed Mixture of Adapter Experts (MoAE) module is integrated into the ViT layers to create MoAE layers during joint fine-tuning. MLPI and MLPV denote the classification heads for SFER and DFER, while FFN, Norm, and MHSA represent the feed-forward network, layer normalization, and multi-head self-attention mechanisms, respectively.
II-A Dynamic Facial Expression Recognition
The evolution of Dynamic Facial Expression Recognition (DFER) has undergone a significant paradigm shift, transitioning from traditional laboratory-based approaches relying on handcrafted features to advanced deep learning-based methods capable of handling in-the-wild scenarios. Recent DFER approaches can be broadly categorized into four distinct groups. The first group, end-to-end supervised learning, initially employed 3D CNNs [22] such as C3D [23], R(2+1)D [24], and I3D-RGB [25] to learn spatiotemporal features from raw videos. This approach evolved to a two-stage process, combining 2D CNNs [22] for spatial feature extraction with RNNs RNNs [26] or LSTMs [27, 10, 28] for temporal modeling. More recently, Transformer-based architectures [29], such as those explored in works like[30, 31, 32], have become dominant in this category. The second group leverages advanced vision-language models, particularly CLIP [33], to advance DFER techniques. Prominent methods such as CLIPER [34], DFER-CLIP [5], and A3lign-DFER [35] capitalize on CLIP’s semantic understanding, enhancing expression recognition by bridging visual and textual modalities. The third group employs self-supervised learning to exploit large amounts of unlabeled facial video data [6, 36, 37]. Notably, MAE-DFER [6] utilizes a local-global interaction Transformer encoder for masked reconstruction, improving task-specific learning. The fourth and final group explores transfer learning from static knowledge to enhance DFER performance [7, 16]. For instance, AEN [16] combines multi-level semantic features from SFER models with emotion-guided loss functions, while S2D [7] adapts a pre-trained SFER model to DFER via Temporal Modeling Adapters. Unlike these methods, which focus on single-modal or task-specific learning, our S4D framework integrates both static and dynamic facial expression data through a unified dual-modal learning approach, offering a more comprehensive and promising solution for DFER.
II-B Multi-Modal Learning and Unified Modeling
Multi-modal learning has significantly advanced computer vision by integrating diverse data modalities, such as image, text, video, and audio [38, 39, 40, 41]. These techniques have also been applied to FER tasks [34, 35, 37]. Traditional methods often train different modalities independently before aligning them, which may overlook the potential benefits of a unified multi-modal learning approach. Recent advancements have aimed to address this issue by unifying the learning process [42, 43]. For instance, BEVT [44] and OmniMAE [43] utilize a single encoder to handle both image and video modalities during pre-training while CoVER [45] employs a unified model for multiple visual datasets and tasks during fine-tuning. However, these methods often fail to maintain learning consistency between the pre-training and fine-tuning stages. In contrast, S4D introduces a unified learning framework integrating facial images and video processing, ensuring consistent multi-modal learning across stages. This novel approach not only overcomes the limitations of previous methods but also significantly improves DFER performance.
III Methodology
III-A Overview
Fig. 2 illustrates the architecture of the S4D framework, which consists of two main stages: Dual-Modal Pre-Training and Joint Fine-Tuning. In the pre-training phase, the framework utilizes the Vision Transformer (ViT) [17] with Masked Autoencoder (MAE) [18] to reconstruct both facial images and videos, learning powerful spatiotemporal representations. After pre-training, the ViT encoder is jointly fine-tuned on both static and dynamic FER datasets in a multi-task learning setup. The final representations, obtained from either images or videos, are fed into separate classification heads for SFER and DFER tasks, with the network jointly optimized via cross-entropy loss. To address the potential risk of negative transfer during direct multi-task joint fine-tuning, the Mixture of Adapter Experts (MoAE) modules are introduced in the deeper layers of ViT to enable generic and task-specific knowledge acquisition. The Dual-Modal Pre-Training, Joint Fine-Tuning, and the MoAE module will be detailed in the following Section III-B, Section III-C, and Section III-D, respectively.
III-B Stage 1: Dual-Modal Self-Supervised Pre-Training
During the pre-training phase, we employ the MAE [21] strategy to jointly train a standard ViT model on large-scale facial image and video datasets. MAE is a self-supervised learning approach that randomly masks a portion of the input data and trains the model to reconstruct the original data from the masked input. This encourages the model to learn powerful representations by capturing the underlying structure and semantics of the data.
Following OmniMAE [43], we treat both image and video inputs as 4D tensors $\bm{X}\in\mathbb{R}^{T\times H\times W\times C}$, where $T$, $H$, $W$, and $C$ represent the number of frames, height, width, and channels, respectively. In this context, we consider an image as a special case of a video with a single frame, thus setting $T=1$. Given an input tensor $\bm{X}$ sampled from image or video datasets, we first generate a random binary mask tensor $\bm{M}^{T\times H\times W}\in\{0,1\}$ with a predefined mask ratio to decide where to drop the patches. The masked input tensor $\bm{X}_{m}$ is computed as the element-wise product of $\bm{X}$ and $\bm{M}$:
$\bm{X}_{m}=\bm{X}\odot\bm{M}.$ (1)
$\bm{X}_{m}$ is then fed into the ViT encoder $f_{E}$ to obtain the latent representation $\bm{Z}=f_{E}(\bm{X}_{m})$. Subsequently, the decoder $f_{D}$ takes $\bm{Z}$ as input and attempts to reconstruct the masked pixels of $\bm{X}$. The objective of the pre-training phase is to minimize the reconstruction loss, which is defined as the mean squared error (MSE) between the reconstructed tensor $\hat{\bm{X}}=f_{D}(\bm{Z})$ and the unmasked tensor $\bm{X}$:
$\mathcal{L}_{\mathrm{MASK}}=\frac{1}{\sum\bm{M}}\sum(1-\bm{M})\odot(\bm{\hat{X}}-\bm{X})^{2}.$ (2)
III-C Stage 2: Joint Fine-Tuning on Static and Dynamic Data
We propose a joint fine-tuning strategy utilizing both SFER and DFER datasets to maintain consistency with dual-modal pre-training and fully exploit knowledge from both domains. This approach mines complementary spatial information from SFER data and temporal dynamics from DFER data, enabling the model to learn more robust, generalizable representations. By integrating these two information sources, the model gains a deeper and more comprehensive understanding of dynamic facial expressions.
As illustrated in Fig. 2, the pre-trained ViT encoder, $f_{E}$, is employed to initialize our S4D encoder, $f_{U}$. The S4D encoder is then jointly fine-tuned on SFER and DFER tasks using the provided inputs ${(\bm{X}_{i},y_{i})}$, where $\bm{X}_{i}$ represents the visual inputs and $y_{i}$ denotes the corresponding labels. During the fine-tuning process, $f_{U}$ generates a unified embedding $\bm{\Phi}=f_{U}(\bm{X})$ for both image and video inputs. The final prediction for each task is generated by a separate task-specific Multi-Layer Perceptron (MLPI for SFER and MLPV for DFER) applied to the final representation $\bm{\Phi}$. To optimize the model, we minimize the cross-entropy loss on the training datasets using mini-batch stochastic gradient descent. Each mini-batch is constructed independently from the SFER or DFER datasets. This approach employs single-source mini-batch sampling, maximizing GPU efficiency by eliminating the need to pad token sequences due to differences in the number of patches between images and videos.
The total loss for joint fine-tuning is defined as follows:
$\mathcal{L}_{\mathrm{total}}=(1-\alpha)\cdot\mathcal{L}_{\mathrm{SFER}}+\alpha\cdot\mathcal{L}_{\mathrm{DFER}},$ (3)
where $\mathcal{L}_{\mathrm{SFER}}$ and $\mathcal{L}_{\mathrm{DFER}}$ are the cross-entropy losses for SFER and DFER tasks, respectively. The binary indicator $\alpha\in\{0,1\}$ toggles between SFER ($\alpha=0$) and DFER ($\alpha=1$) tasks based on the mini-batch source.
III-D Mixture of Adapter Experts Module
During the fine-tuning stage, we incorporate static expression data alongside DFER datasets to improve the model’s performance. However, using simple multi-task learning approaches, such as multiple classification heads, may lead to negative transfer, failing to fully exploit the correlation between SFER and DFER, thus hindering optimal performance. To address this issue and fully utilize the correlations between different SFER and DFER tasks, we propose the Mixture of Adapter Experts (MoAE) module.
The design of the MoAE module is inspired by the Mixture of Experts (MoE) [46], which employs multiple expert networks coordinated by a gating mechanism. In the MoE framework, a gating network $G$ is responsible for assigning weights to $n$ independent experts $\{E_{i}\}^{n}_{i=1}$ based on the input $\bm{x}$. The gating network computes a weight distribution by applying a $\operatorname{Softmax}$ function to the dot product of the input and a learnable matrix $\bm{W}_{g}$:
$G(\bm{x})=\operatorname{Softmax}(\bm{x}\cdot{\bm{W}_{g}}).$ (4)
To encourage load balance, we adopt Noisy Top-K Gating [47]. This gating mechanism introduces noise to the logits before applying the $\operatorname{Top-K}$ operation and $\operatorname{Softmax}$ function:
$H(\bm{x})=\bm{x}\cdot\bm{W}_{g}+\epsilon,$ (5)
$G(\bm{x})=\operatorname{Softmax}(\operatorname{Top-K}(H(\bm{x}),k)),$ (6)
where $\epsilon\sim\mathcal{N}(0,\sigma^{2}\bm{I})$ is Gaussian noise with mean 0 and variance $\sigma^{2}$, and $\operatorname{Top-K}(H(\bm{x}),k)$ retains the top $k$ largest values from $H(\bm{x})$ while suppressing the remaining elements to negative infinity. The introduction of noise helps to diversify the expert selection, preventing the gating network from always choosing the same experts. The output of the MoE layer is then computed as a weighted sum of the expert outputs:
$\operatorname{MoE}(\bm{x})=\sum_{i=1}^{n}G(\bm{x})_{i}\cdot E_{i}(\bm{x}).$ (7)
Previous works, such as ViT-MoE [48], often use Feed-Forward Networks (FFNs) as experts in the MoE, significantly increasing the model parameters but potentially compromising structure and performance. To preserve the model structure and learn task-specific knowledge with minimal parameter increase, we employ parameter-efficient adapters [49] as experts in MoAE. These adapters are lightweight modules consisting of two linear layers with a $\operatorname{GELU}$ activation function:
$\operatorname{Adapter}(\bm{x})=\bm{x}+\bm{W}_{2}(\operatorname{GELU}(\bm{W}_{1}\bm{x}+\bm{b}_{1}))+\bm{b}_{2},$ (8)
where $\bm{W}_{1}\in\mathbb{R}^{d\times r},\bm{W}_{2}\in\mathbb{R}^{r\times d},\bm{b}_{1}\in\mathbb{R}^{r},\bm{b}_{2}\in\mathbb{R}^{d}$ are learnable parameters. The input/output dimension $d$ and bottleneck dimension $r$ (set to $d/4$ in our work) satisfy $r\ll d$, allowing the adapter to learn representations with minimal additional parameters. The formulation of the MoAE module can be expressed as:
$\operatorname{MoAE}(\bm{x})=\sum_{i=1}^{n}G(\bm{x})_{i}\cdot\operatorname{Adapter}_{i}(\bm{x}),$ (9)
where $\operatorname{Adapter}_{i}$ denotes the $i$-th adapter expert and $G(\bm{x})$ is the gating network determining each adapter expert’s contribution based on input $\bm{x}$.
Unlike ViT-MoE, which replaces the FFN with MoE, we integrate the MoAE module into the ViT layer, resulting in the MoAE layer, which operates in parallel with the original FFN. This design, as shown in Fig. 2, replaces the latter half of the ViT layers with MoAE layers, preserving the original structure while enabling the FFN to focus on task-agnostic knowledge and the MoAE to capture task-specific insights. Additionally, our approach differs from the Mixture of Parameter-Efficient Experts method [50] commonly used in large language models by mitigating negative transfer and enhancing adaptability. The computation flow in the MoAE layer is as follows:
$\bm{x}^{\prime}=\bm{x}+\operatorname{MHSA}(\operatorname{LayerNorm}(\bm{x})),$ (10)
where $\operatorname{MHSA}$ is multi-head self-attention, $\operatorname{LayerNorm}$ is layer normalization, and $\bm{x}^{\prime}$ represent the global relational representation. $\bm{x}^{\prime}$ is then processed by FFN and MoAE:
$\bm{x}_{g}=\operatorname{FFN}(\operatorname{LayerNorm}(\bm{x}^{\prime})),$ (11)
$\bm{x}_{s}=\operatorname{MoAE}(\operatorname{LayerNorm}(\bm{x}^{\prime})),$ (12)
where $\bm{x}_{g}$ and $\bm{x}_{s}$ represent generic knowledge and task-specific knowledge, respectively. Finally, the output $\bm{x}_{o}$ of MoAE layer is calculated as:
$\bm{x}_{o}=\bm{x}^{\prime}+\bm{x}_{g}+\bm{x}_{s}.$ (13)
IV Experiments
IV-A Datasets
Pre-Training Datasets. We conduct dual-modal self-supervised pre-training on two large-scale facial datasets: VoxCeleb2 [12] and AffectNet [51]. VoxCeleb2 is a comprehensive audio-visual speaker recognition dataset sourced from YouTube, comprising over 1 million utterances by 6,112 speakers. For pre-training, we utilize the development set, which includes 1,092,009 video clips extracted from 145,569 videos. AffectNet is a large-scale static facial expression dataset containing over 1 million images, of which 450,000 are manually annotated. AffectNet is divided into two subsets: AffectNet-8, featuring eight basic expressions, and AffectNet-7, featuring seven basic expressions. A total of 287,568 images from the AffectNet-8 training set are combined with the VoxCeleb2 dataset for the pre-training process.
Fine-Tuning Datasets: We evaluate our method on three widely recognized DFER benchmarks: DFEW [11], FERV39K [9], and MAFW [10]. The DFEW dataset comprises 16,372 video clips sourced from over 1,500 films, each annotated with one of seven basic expressions. FERV39K, the largest in-the-wild DFER dataset, includes 38,935 clips representing seven basic expressions across diverse scenarios, divided into 31,088 training clips and 7,847 testing clips. MAFW is a multimodal DFER dataset containing 10,045 video clips; however, our focus is solely on the video modality, which consists of 9,172 clips categorized into 11 classes. For the MAFW dataset, we employed MTCNN [52] to extract and align facial regions, effectively removing irrelevant backgrounds. Additionally, AffectNet-7 [51] was utilized for joint training alongside the DFER datasets. Table I summarizes the key characteristics of the above datasets.
TABLE I: A summary of the basic information about the pre-training and fine-tuning datasets used in this paper. V: video, I: image, CV: cross-validation.
Dataset Samples Classes Modality Evaluation
Pre-Training Datasets
VoxCeleb2 (dev) 1,092,009 - V -
AffectNet-8 (train) 287,568 - I -
Fine-Tuning Datasets
DFEW 11,697 7 V 5-fold CV
FERV39K 38,935 7 V Train & test
MAFW 10,045 11 V 5-fold CV
AffectNet-7 287,401 7 I Train & test
We evaluate the performance using unweighted average recall (UAR) and weighted average recall (WAR), consistent with previous studies [6, 7]. For DFEW and MAFW, we employ 5-fold cross-validation, aggregating predictions and labels across all folds for final UAR and WAR computation.
TABLE II: Comparisons of our S4D with the state-of-the-art DFER methods on DFEW, FERV39k, and MAFW. Baseline results are directly extracted from [6]. The best results are highlighted in bold, and the second-best underlined. TI:Time Interpolation; DS: Dynamic Sampling.
Method Sample Strategies Backbone DFEW FERV39k MAFW
UAR (%) WAR (%) UAR (%) WAR (%) UAR (%) WAR (%)
EC-STFL  [11] TI C3D / P3D 45.35 56.51 - - - -
Former-DFER  [32] DS Transformer 53.69 65.70 37.20 46.85 31.16 43.27
CEFLNet  [53] Clip-based ResNet-18 51.14 65.35 - - - -
NR-DFERNet  [54] DS CNN-Transformer 54.21 68.19 33.99 45.97 - -
STT  [30] DS ResNet-18 54.58 66.65 37.76 48.11 - -
EST  [55] DS ResNet-18 53.94 65.85 - - - -
Freq-HD  [56] FreqHD VGG13-LSTM 46.85 55.68 33.07 45.26 - -
LOGO-Former  [57] DS ResNet-18 54.21 66.98 38.22 48.13 - -
IAL  [31] DS ResNet-18 55.71 69.24 35.82 48.54 - -
AEN  [16] DS ResNet-18 56.66 69.37 38.18 47.88 - -
T-MEP [58] DS Transformer 57.16 68.85 - - 39.37 52.85
M3DFEL  [59] DS ResNet-18-3D 56.10 69.25 35.94 47.67 - -
CLIPER  [34] DS CLIP-ViT-B/16 57.56 70.84 41.23 51.34 - -
DFER-CLIP  [5] DS CLIP-ViT-B/32 59.61 71.25 41.27 51.65 39.89 52.55
EmoCLIP  [60] DS CLIP-ViT-B/32 58.04 62.12 31.41 36.18 34.24 41.46
A3lign-DFER  [35] DS CLIP-ViT-L/14 64.09 74.20 41.87 51.77 42.07 53.24
SVFAP  [36] DS ViT-B/16 62.83 74.27 42.14 52.29 41.19 54.28
HiCMAE  [37] DS ViT-B/16 63.76 75.01 - - 42.65 56.17
MAE-DFER  [6] DS ViT-B/16 63.41 74.43 43.12 52.07 41.62 54.31
S2D  [7] DS ViT-B/16 61.82 76.03 41.28 52.56 41.86 57.37
VideoMAE (Baseline)  [6] DS ViT-B/16 63.60 74.60 43.33 52.39 40.87 53.51
S4D (Ours) DS ViT-B/16 66.80 76.68 43.41 53.65 43.72 58.44
IV-B Implementation Details
We adopt ViT-B/16 [17] as the backbone for S4D and VideoMAE [8, 6] as the baseline. To reduce background noise, we crop a central 160$\times$160 patch from each 224$\times$224 frame in the VoxCeleb2 dataset, following [6]. For pre-training, we sample 16 frames per clip with a temporal stride of 4, use a patch size of $2\times 16\times 16$. Static images are resized to 160$\times$160 and temporally replicated to match the patch size. We apply random masking at a ratio of 95% for videos and 90% for images, consistent with [43]. Training is performed using the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$, an over batch size $N_{\mathrm{bs}}$ of 384, and a base learning rate $lr_{\mathrm{base}}$ of 1.6e-3, with weight decay set to 0.05. The learning rate is scaled linearly based on batch size: $lr=lr_{\mathrm{base}}\times\frac{N_{\mathrm{bs}}}{512}$. Pre-training is conducted on the VoxCeleb2 and AffectNet-8 datasets for 100 epochs, with a cosine learning rate scheduler.
For fine-tuning, we replace the latter half of the ViT layers with MoAE layers and initialize the model using the pre-trained ViT encoder weights. All input frames and images are resized to $160\times 160$. The batch size is 32 for the DFER datasets and 64 for the SFER datasets, with a learning rate of $4\times 10^{-5}$. We sample 16 frames per clip with a temporal stride of 4 across all datasets, except for FERV39K, which uses a stride of 1 due to its unique properties. In the MoAE module, we set $k=2$ and the number of experts $n=8$. During joint fine-tuning, the proportion of the SFER dataset in each epoch is empirically set to 50%. We optimize the model using AdamW for 100 epochs. During inference, following [6, 7], we uniformly sample two clips per video and averaging their predictions for DFER tasks. Unless otherwise specified, all results are obtained using the S4D (ViT+MoAE) model with the best-performing weights.
All pre-training experiments are performed on two Nvidia A800 GPUs, and fine-tuning experiments on two 4090 GPUs, utilizing the PyTorch framework.
TABLE III: Comparative analyses of accuracy across various emotion categories: S4D vs. other approaches on DFEW. The best results are highlighted in bold, and the second-best underlined. Baseline results are directly extracted from [6].
Method Accuracy of Each Emotion DFEW
Happy Sad Neutral Angry Surprise Disgust Fear UAR (%) WAR (%)
C3D  [23] 75.17 39.49 55.11 62.49 45.00 1.38 20.51 42.74 53.54
R(2+1)D-18  [24] 79.67 39.07 57.66 50.39 48.26 3.45 21.06 42.79 53.22
3D ResNet-18  [61] 76.32 50.21 64.18 62.85 47.52 0.00 24.56 46.52 58.27
EC-STFL  [11] 79.18 49.05 57.85 60.98 46.15 2.76 21.51 45.35 56.51
ResNet-18+LSTM  [32] 83.56 61.56 68.27 65.29 51.26 0.00 29.34 51.32 63.85
ResNet-18+GRU  [32] 82.87 63.83 65.06 68.51 52.00 0.86 30.14 51.68 64.02
Former-DFER  [32] 84.05 62.57 67.52 70.03 56.43 3.45 31.78 53.69 65.70
CEFLNet [53] 84.00 68.00 67.00 70.00 52.00 0.00 17.00 51.14 65.35
NR-DFERNet  [54] 88.47 64.84 70.03 75.09 61.60 0.00 19.43 54.21 68.19
STT  [30] 87.36 67.90 64.97 71.24 53.10 3.49 34.04 54.58 66.65
EST  [55] 86.87 66.58 67.18 71.84 47.53 5.52 28.49 53.43 65.85
IAL  [31] 87.95 67.21 70.10 76.06 62.22 0.00 36.44 55.71 69.24
M3DFEL  [59] 89.59 68.38 67.88 74.24 59.69 0.00 31.64 56.10 69.25
MAE-DFER  [6] 92.92 77.46 74.56 76.94 60.99 18.62 42.35 63.41 74.43
SVFAP  [36] 93.13 76.98 72.31 77.54 65.42 15.17 39.25 62.83 74.27
S2D  [7] 93.62 80.25 77.14 81.09 64.53 1.38 34.71 61.82 76.03
VideoMAE (Baseline)  [6] 93.09 78.78 71.75 78.74 63.44 17.93 41.46 63.60 74.60
S4D (Ours) 94.76 79.04 74.60 79.80 67.26 27.59 44.57 66.80 76.68
IV-C Comparison with the State of the Art
To evaluate the performance of our S4D method, we compared it with state-of-the-art methods on three publicly available in-the-wild DFER datasets: DFEW [11], FERV39K [9], and MAFW [10]. The results are summarized in Table II.
As shown in Table II, S4D achieves the best performance across all three datasets, significantly outperforming previous state-of-the-art methods, including A3lign-DFER [35], MAE-DFER [6], and S2D [7]. Specifically, S4D outperforms S2D by 0.65% WAR on DFEW, 1.09% WAR on FERV39k, and 1.07% WAR on MAFW. Additionally, S4D shows substantial improvements in UAR, outperforming S2D by 4.98% on DFEW, 2.13% on FERV39k, and 1.92% on MAFW. These improvements highlight that our method effectively learns robust representations and mitigates the impact of class imbalance. Compared to the baseline method VideoMAE [6], S4D achieves significant improvements of +3.20%/+2.08% UAR/WAR on DFEW, +0.08%/+1.26% UAR/WAR on FERV39k, and +2.87%/+4.93% UAR/WAR on MAFW. These results demonstrate S4D’s effectiveness in learning representations through dual-modal pre-training and joint fine-tuning on both static and dynamic FER data, underscoring its superiority in DFER tasks.
Additionally, Table III presents a detailed analysis of S4D’s category-specific performance on the DFEW dataset, demonstrating its strong overall accuracy. Specifically, S4D achieves the highest accuracy in happy (94.76%), surprise (67.26%), disgust (27.59%), and fear (44.57%), with a notable 9.66% improvement in disgust over the baseline method VideoMAE [6]. It also ranks second in sad (79.04%), neutral (74.60%), and angry (79.80%). Notably, S4D excels in traditionally challenging categories (surprise, fear, and disgust), where data imbalance is prevalent, due to its effective dual-modal pre-training and joint fine-tuning strategy. This approach enables S4D to construct representations with reduced bias while effectively leveraging the complementary information from SFER data through unified dual-modal learning, resulting in more robust emotion recognition across all categories. Although the improvement in disgust accuracy is noteworthy, the relatively low absolute accuracy (27.59%) underscores the inherent difficulty in recognizing this subtle emotion and suggests potential data imbalance within the DFEW dataset. Further investigation into techniques for mitigating data imbalance may yield additional improvements in this category.
IV-D Ablation Studies
To evaluate the effectiveness of the key components in the S4D framework, we conduct ablation studies on the FERV39K [9], DFEW [11], and AffectNet-7 [51] datasets. For simplicity and to reduce computational costs, we report the results for DFEW based on one fold of the 5-fold cross-validation.
IV-D1 Impact of Dual-Modal Pre-training
Table IV analyzes the impact of various pre-training settings on DFER performance. All pre-training methods result in significant improvements in both UAR and WAR compared to the no pre-training setting, emphasizing the importance of pre-training in DFER. Specifically, pre-training on single-modal video data outperforms image-based pre-training, as it captures crucial temporal information. However, image-based pre-training also proves beneficial by enabling the model to learn fine-grained facial features that complement the temporal representations derived from video data. Dual-modal pre-training, which combines image (AffectNet) and video (VoxCeleb2) datasets (Experiment #4), achieves the best performance, with a UAR/WAR of 43.41%/53.65% on FERV39K and 70.58%/79.37% UAR/WAR on DFEW. This confirms that dual-modal pre-training effectively captures both temporal information and fine-grained facial details, leading to robust representation capabilities. Experiment #5, which utilized homogeneous image and video data from VoxCeleb2 for dual-modal pre-training, exhibited slightly lower performance than Experiment #4. This suggests that leveraging diverse datasets and modalities for pre-training can yield richer and more beneficial feature representations. Overall, dual-modal pre-training significantly enhances DFER performance, surpassing both single-modal and no pre-training approaches.
TABLE IV: Comparison of the proposed dual-modal pre-training with other pre-training settings. Dual-modal pre-training on AffectNet and VoxCeleb2 significantly improves DFER performance and outperforms single-modal pre-training settings with less pre-training cost. AFC: AffectNet; VC2: VoxCeleb2.
# Pre-Training Setting (images + videos) Pre-Training Cost (TFLOPs) FERV39K DFEW
UAR WAR UAR WAR
1 Without pre-training 0 27.67 39.56 40.96 52.65
2 Image only (AFC) $8.8\times 10^{4}$ 36.10 47.59 53.80 66.82
3 Video only (VC2) $1.8\times 10^{6}$ 42.92 52.58 67.71 76.33
4 Dual-modal (AFC + VC2) $\bm{1.4\times 10^{6}}$ 43.41 53.65 70.58 79.37
5 Dual-modal (VC2 + VC2) $1.4\times 10^{6}$ 42.36 52.94 68.35 78.72
IV-D2 Pre-Training Cost Analysis
As shown in Table IV, the proposed dual-modal pre-training (Experiment #4 and #5) demonstrates superior efficiency and effectiveness, significantly reducing computational costs compared to video-only pre-training (Experiment #3). Specifically, the integration of image data reduces the pre-training cost by approximately 22% (from 1.8 to 1.4 $\times$ 106 TFLOPs), achieved through aggressive masking ratios of 90% for images and 95% for videos, minimizing redundant computations. This approach not only optimizes efficiency but also enhances performance, yielding notable improvements on the FERV39K (+0.49% UAR, +1.07% WAR) and DFEW (+2.87% UAR, +3.04% WAR) benchmarks. In contrast, image-only pre-training (Experiment #2) incurs a minimal cost (8.8 $\times$ 104 TFLOPs) but shows limited effectiveness for DFER tasks. Similarly, the no pre-training setting (Experiment #1) incurs no pre-training cost but severely degrades performance. These findings underscore the critical importance of capturing temporal dynamics for DFER, which are absent in image-only pre-training or no pre-training strategies.
TABLE V: Performance and parameter efficiency of S4D compared to other multi-task learning (MTL) methods. ViT-MTL refers to the direct application of ViT in MTL tasks with multiple heads. ViT-MoE [48] replaces the FFN in the ViT layer with an MoE and employs FFNs as experts.
# Model Setting Params (M) FERV39K DFEW
Test / Train UAR WAR UAR WAR
6 ViT 86 / 86 42.68 53.08 68.69 77.01
7 ViT-MTL (w/o MoAE) 86 / 86 42.63 52.96 69.62 78.08
8 S4D (Ours) 90 / 101 43.41 53.65 70.58 79.37
9 ViT-MoE 115 / 285 41.79 52.81 68.88 78.12
IV-D3 Evaluation of the MoAE Module
To evaluate MoAE’s effectiveness in mitigating negative transfer between SFER and DFER tasks, we conducted comprehensive comparisons with other multi-task learning (MTL) methods. As shown in Table V, the vanilla ViT model achieves a performance of 42.68%/53.08% UAR/WAR on FERV39K and 68.69%/77.01% UAR/WAR on DFEW dataset. When SFER data is incorporated through simple multi-task learning (simply adding multiple task-specific heads to ViT, Experiment #7), we observe minimal improvement or even slight performance degradation compared to the vanilla ViT. This phenomenon suggests that straightforward multi-task adaptation may fail to capture task relationships and lead to negative transfer. In contrast, our proposed S4D (with MoAE) demonstrates significant improvements across all metrics while maintaining reasonable parameter efficiency. Specifically, S4D achieves the best performance with 43.41%/53.65% UAR/WAR on FERV39K and 70.58%/79.37% UAR/WAR on DFEW, validating its effectiveness in learning both task-specific knowledge and generic features. To further investigate the architectural benefits, we compared S4D with ViT-MoE [48], a classic MoE architecture that replaces the standard FFN with FFN experts. Despite utilizing substantially more parameters (115M/285M vs. 90M/101M for testing/training), ViT-MoE exhibits poorer performance. These results emphasize that S4D’s superior performance stems from its architectural innovations rather than mere parameter scaling, highlighting its parameter efficiency and effectiveness in learning complementary knowledge between SFER and DFER tasks while mitigating negative transfer.
TABLE VI: Ablation studies on the joint fine-tuning strategy. The joint fine-tuning strategy significantly enhances performance, demonstrating the complementary benefits of integrating static and dynamic expression data. *: Produced by the epoch at 43 for optimal SFER performance.
Fine-Tuning Setting (data mixture) FERV39K AffectNet-7
UAR WAR UAR WAR
SFER (Static) - - 65.58 65.64
DFER (Dynamic) 42.45 52.82 - -
Joint (50% Static + 100% Dynamic) 43.41 53.65 65.45 65.50
Joint (50% Static + 100% Dynamic) 40.52 50.90 65.88 65.94
IV-D4 The Effectiveness of Joint Fine-Tuning
We conducted an ablation study to assess the effectiveness of joint fine-tuning on both SFER and DFER tasks, as presented in Table VI. Our findings reveal that joint fine-tuning significantly outperforms individual fine-tuning, particularly for DFER. On the FERV39K dataset, joint fine-tuning improves UAR/WAR by 0.96%/0.83% over DFER-only fine-tuning. For SFER on the AffectNet-7 dataset, it improves UAR/WAR by 0.30%/0.30%. These results underscore the complementary nature of static and dynamic expression knowledge and demonstrate the effectiveness of joint fine-tuning in enhancing expression representation learning. They also highlight S4D’s ability to handle both SFER and DFER tasks simultaneously, effectively leveraging the strengths of each modality. However, achieving optimal results required different training epochs for DFER and SFER, which indicates challenges in simultaneous optimization.
TABLE VII: Contribution of proposed components. D.P.: Dual-Modal Pre-training, J.F.: Joint Fine-tuning.
D.P. MoAE J.F. FERV39K DFEW
UAR WAR UAR WAR
42.58 52.34 64.66 76.16
42.68 53.08 68.69 77.01
42.45 52.82 66.52 76.93
43.41 53.65 70.58 79.37
IV-D5 Contribution of Proposed Components
We conducted ablation studies to evaluate the impact of each component in our S4D framework: dual-modal pre-training (D.P.), MoAE, and joint fine-tuning (J.F.). Results from Table VII indicate that D.P. improved UAR/WAR by 0.10%/0.74% on FERV39K and 4.03%/0.85% on DFEW over the baseline (our fine-tuned version of VideoMAE on the pre-trained weights from [6]), underscoring its role in enhancing spatiotemporal representations. However, MoAE alone led to a slight performance decrease on all metrics, suggesting it may require a joint fine-tuning strategy to unlock its task-specific capabilities fully. The optimal configuration, integrating all components, achieved a UAR/WAR of 43.41%/53.65% on FERV39K and 70.58%/79.37% on DFEW datasets, highlighting the synergistic benefits of our S4D framework in refining DFER performance through discriminative feature learning.
IV-D6 Ablation Study on the Position of MoAE
To identify the optimal placement of the MoAE module within the ViT architecture, we conducted ablation studies at different positions. Table VIII shows that placing the MoAE in the last six layers (Later) yields the highest UAR/WAR performance at 43.41%/53.65% on FERV39K and 70.58%/79.37% on DFEW, respectively. In contrast, placing the MoAE in the first six layers (Early) results in a lower UAR/WAR of 43.15%/53.46% on FERV39K and 67.93%/78.42% on DFEW, suggesting that early placement may hinder the model’s ability to learn generic features. Middle and alternate layer placements produce intermediate results. These findings indicate that the MoAE module is most effective in the later layers, where it can enhance task-specific learning by leveraging the learned generic representations.
TABLE VIII: Ablation studies on the position of the MoAE module within the ViT architecture.
Position FERV39K DFEW
UAR WAR UAR WAR
Early 43.15 53.46 67.93 78.42
Middle 43.32 53.57 69.96 79.15
Later 43.41 53.65 70.58 79.37
Alternate 43.27 53.52 70.48 79.20
IV-D7 Ablation Studies on Hyperparameters
We conducted comprehensive ablation studies to investigate the impact of three critical hyperparameters: the number of experts ($n$) in the MoAE module, the number of MoAE layers, and the proportion of the SFER dataset utilized during dual-modal pre-training and joint fine-tuning.
Our experiments first focused on the structural parameters of the MoAE module. The results indicate that the model achieves optimal performance in both the FERV39K and DFEW datasets when the MoAE module is configured with eight experts, as illustrated in Fig. 3(a). Any deviation from this optimal number, whether an increase or decrease, results in marginal performance degradation. Similarly, our investigation into the impact of MoAE layer count reveals that incorporating six MoAE layers yields the best performance, as shown in Fig. 3(b). Deviating from this number leads to a noticeable decline in performance. These findings suggest that replacing the final six ViT layers with MoAE layers establishes an optimal balance between task-specific adaptation and preservation of general knowledge.
Beyond structural parameters, we also explored the impact of external training data. As depicted in Fig. 3(c), increasing the proportion of the SFER dataset during the pre-training phase leads to consistent performance improvements on the DFEW dataset (up to 2.87%/3.04% UAR/WAR gain), whereas the impact on FERV39k is less pronounced (approximately 0.49%/1.07% UAR/WAR improvement). These differences may reflect variations in dataset characteristics, such as scale, label noise, or inherent task difficulty. Both datasets achieve peak performance when the entire SFER dataset is employed during pre-training, indicating that pre-training on SFER data is generally advantageous, especially for DFEW. Regarding the SFER dataset utilization during fine-tuning, our experiments reveal that employing 50% of the SFER data per epoch produces optimal results, as depicted in Fig. 3(d). However, increasing the proportion beyond 50% offers no additional benefits and may potentially bias the optimization trajectory of the DFER task.
Overall, these empirical findings highlight the importance of careful hyperparameter tuning in maximizing the effectiveness of the proposed unified learning framework. Specifically, using eight MoAE experts, six MoAE layers, and the entire SFER dataset during pre-training, along with partial SFER data during fine-tuning, establishes a robust foundation for practical deployment while maintaining computational efficiency.
(a) Total number of experts $n$.
(b) The number of MoAE layers.
(c) Proportion of SFER dataset used during pre-training.
(d) Proportion of SFER dataset used during fine-tuning.
Figure 3: Analyses of the total number of experts, the number of MoAE layers, and the proportion of SFER data used during dual-modal pre-training and joint fine-tuning.
IV-E Correlation Analysis between SFER and DFER
IV-E1 Evaluation on Cross-Task
Since S4D can handle both SFER and DFER tasks using a shared backbone and respective classifiers, we treat it as a distinct task model during inference. To investigate the correlations and differences between the SFER and DFER tasks, we conducted cross-task evaluation comparisons on the AffectNet-7, FERV39K, and DFEW datasets. As shown in Table IX, the SFER model trained on AffectNet-7 achieves 32.34%/35.10% and 43.36%/47.13% UAR/WAR on the FERV39K and DFEW datasets, respectively. Although this performance is lower than that of DFER models on their respective tasks, it significantly exceeds random guessing (approximately 14.3%), indicating that the SFER model captures relevant spatial features from video data. In contrast, DFER models trained on the FERV39K and DFEW datasets achieve 51.46%/51.38% and 45.20%/45.04% UAR/WAR on AffectNet-7, respectively. These results suggest that DFER models can leverage static expression information from dynamic data, transferring well to the SFER task. We attribute this phenomenon to the inherent presence of static facial expressions within dynamic sequences, which capture peak expressions and serve as snapshots within the temporal progression of the video. However, direct comparisons of UAR/WAR metrics between SFER and DFER models are not ideal due to the inherent differences between the tasks. Therefore, these results should be viewed as reflecting feature transferability rather than a direct performance comparison.
TABLE IX: Cross-task evaluation comparisons on SFER (AffectNet-7) and DFER (FERV39K, DFEW).
Train Test UAR WAR
AffectNet-7 FERV39K 32.34 35.10
AffectNet-7 DFEW 43.36 47.13
FERV39K AffectNet-7 51.46 51.38
DFEW AffectNet-7 45.20 45.04
IV-E2 Semantic Relevance between SFER and DFER Tasks
Figure 4: The semantic relevance between SFER and DFER tasks. NE, HA, SA, AN, SU, DI, and FE denote neutral, happy, sad, anger, surprise, disgust, and fear, respectively.
To further investigate the semantic relevance between SFER and DFER tasks, we analyzed the cosine similarity between the class representation centers of FERV39K and AffectNet-7, as well as between DFEW and AffectNet-7. As illustrated in Fig. 4, the similarity matrix reveals prominently high values along the diagonal, indicating strong semantic correlations between corresponding expression classes across the two tasks. These findings provide robust empirical evidence supporting our motivation to utilize SFER data to enhance the understanding of dynamic expressions and improve DFER performance. Moreover, additional regions of elevated similarity were observed, particularly related to the neutral expression category, suggesting inherent semantic overlaps among certain emotion categories.
(a) FERV39K vs. AffectNet-7.
(b) DFEW vs. AffectNet-7.
Figure 5: Visualization of the activation distribution of experts in MoAE layers: FERV39K vs. AffectNet-7 and DFEW vs. AffectNet-7 datasets.
Figure 6: Visualizations of original frames (first row) and attention maps (only positive values) from baseline model (second row) and S4D (third row) across four emotion categories: sad, fear, angry, surprise (from left to right, top to bottom).
(a) DFEW
(b) FERV39K
(c) MAFW
Figure 7: Confusion matrices of S4D on DFEW, FERV39K and MAFW datasets. AN: angry. DI: disgust. FE: fear. HA: happy. NE: neutral. SA: sad. SU: surprise. CO: contempt. AX: anxiety. HL: helplessness. DS: disappointment.
IV-E3 Distribution of the Expert Activation in MoAE Layers
Figs. 5(a) and 5(b) show the various activation patterns of experts within the MoAE layers: FERV39K (DFER) vs. AffectNet-7 (SFER) and DFEW (DFER) vs. AffectNet-7 (SFER) datasets. In the FERV39K dataset, expert E3 demonstrates peak activation in layer L4, indicating its specialized capability in processing dynamic facial expressions. Conversely, for AffectNet-7, expert E7 shows pronounced activation in layer L5, suggesting its particular aptitude for SFER. This pattern of task-specific specialization is further validated by the activation patterns observed between DFEW and AffectNet-7 datasets. Specifically, expert E3 exhibits strong activation in layer L4 when processing DFEW data while maintaining relatively lower activation levels for AffectNet-7. Similarly, expert E6 displays higher activation in layer L5 for AffectNet-7, with comparatively reduced activation levels for DFEW. Notably, experts like E1 and E2 display moderate activation across multiple layers in both datasets, suggesting their role in capturing common features shared between tasks. These patterns reveal that while MoAE captures shared representational knowledge between SFER and DFER tasks, it effectively distinguishes task-specific characteristics through expert specialization across layers. Additionally, the similar expert activation patterns observed across different experiments for the AffectNet-7 dataset can be attributed to the same initialization conditions, but also hint at the potential formation of specialized processing pathways for SFER tasks.
IV-F Visualization
IV-F1 Visualization of Attentions
Fig. 6 presents the attention maps from both the baseline model and S4D across four emotion categories: sad, fear, angry, and surprise. The results demonstrate that integrating SFER knowledge enables our model to achieve more focused and precise attention compared to the baseline. Specifically, for the sad expression, the baseline model exhibits relatively dispersed attention, particularly around the mouth and forehead—regions that are generally less reliable for distinguishing sadness. In contrast, the S4D model effectively concentrates on the eyes, brows, and mouth, which are critical regions for recognizing sadness, as indicated by the activation of Action Units (AUs) 1 (Inner Brow Raiser), 4 (Brow Lowerer), and 15 (Lip Corner Depressor) [62]. Similarly, for fear, the baseline model primarily allocates attention to the upper face while neglecting key regions. Our model, however, effectively targets the eyes and mouth, which are essential for detecting fear, aligning with the activation of AUs such as 1 (Inner Brow Raiser), 5 (Upper Lid Raiser), and 26 (Jaw Drop) [62]. This trend is also evident in the attention maps for both angry and surprise expressions, where the S4D model demonstrates a more comprehensive and focused attention allocation than the baseline. In summary, incorporating SFER data significantly enhances the model’s attention to critical facial regions across all emotional categories, leading to a notable improvement in DFER performance over the baseline.
IV-F2 Confusion Matrix Analysis
We performed a comprehensive statistical analysis of S4D’s performance on the DFEW [11], FERV39k [9], and MAFW [10] datasets, with the results presented through confusion matrices. As depicted in Fig. 7, major categories such as happy, sad, and neutral expressions demonstrate high accuracy across all datasets. In contrast, minor categories, including disgust, fear, and helplessness exhibit lower accuracy, particularly on the DFEW dataset. This discrepancy likely arises from the limited sample sizes and data imbalance within these categories. Such challenges can be alleviated by implementing imbalanced learning strategies, such as oversampling techniques. Additionally, we observed that some expressions are frequently misclassified as neutral, particularly disgust and surprise, due to the subtle boundaries between neutral and these emotional expressions. For example, in both the DFEW and FERV39K datasets, disgust is frequently misclassified as neutral, likely because of the subtlety of disgust expressions, such as slight nose wrinkling or minor lip movements.
V Conclusion
In this paper, we proposed Static-for-Dynamic (S4D), a novel unified dual-modal learning framework that integrates SFER data to enhance DFER. By employing dual-modal pre-training and joint fine-tuning on both FER image and video datasets, S4D effectively learns powerful spatiotemporal representations, enabling a more comprehensive understanding of dynamic facial expressions and significantly advancing dynamic expression recognition. Furthermore, the proposed MoAE module, integrated into the latter ViT layers, empowers the model to better utilize learned generic representations for task-specific feature extraction. This design effectively mitigates negative transfer and promotes more discriminative feature learning tailored to DFER. Additionally, despite the differences between SFER and DFER tasks, our analysis of the correlation between these tasks from both semantic and expert pathway perspectives highlights their complementary nature in dual-modal learning. Extensive experimental evaluations on the FERV39K, MAFW, and DFEW benchmarks demonstrated the effectiveness and superiority of S4D, surpassing the baseline by large margins and setting a new state of the art.
However, achieving state-of-the-art performance for DFER requires a careful balance of SFER data during joint fine-tuning, which poses challenges in optimizing both DFER and SFER tasks simultaneously. In future work, we will focus on developing more efficient architectures and optimization strategies to unify the learning of DFER, SFER, and other visual affective recognition tasks, aiming for comprehensive excellence across all aspects of these tasks.